A source emits 2000Hz sound waves. At a distance of 300m from the source, a device is installed that records the arrival of waves through the air and sends them back through a straight metal wire. The length of these waves on the wire is 17m. What is the round trip time of the waves, in seconds?
Given: speed of sound in the air equal to 300m / s.
1
Expert's answer
2020-05-12T10:24:37-0400

"t=t_1+t_2"


in the air


"t_1=\frac{300}{300}=1s"


on the wire


"t_2=\frac{300}{v}" , "v=\lambda\cdot \nu=17\cdot2000=34000m/s"


"t_2=\frac{300}{v}=\frac{300}{34000}\approx0.01s"


the round trip time of the waves


"t=1+0.01=1.01 s".